Leann, with a $300,000 bequest from her father and a business degree from Athabasca University, is considering opening a gift shop in North Edmonton. If her shop is highly successful, she expects an annual net profit of $220,000. If the business is moderately successful, she expects $130,000. If not successful, she expects to have zero net profit. Under any circumstances, she is not contemplating any loss. Her anticipated probabilities of these three options are: 0.5, 0.3 and 0.2, respectively.

a) Calculate her expected net profit. Also calculate the standard deviation of her profit.

b)The business requires a $300,000 investment. If she has a 20% opportunity cost on invested funds of similar riskiness, should the project be undertaken?

c) Suppose Leann considers two alternative investment options instead of opening a gift shop. She has the option to buy a risk free asset that will pay 10%, or she can invest in a stock that has a 0.3 chance of paying 10%, a 0.2 chance of paying 22%, and a 0.5 chance of providing a 20% return. If she invests $160,000 in the stock and $140,000 in the risk free asset, determine the expected percentage return on the stock and the standard deviation.
Answer #1

(a) Making a probability density function table

Profit (x) Probability X * P(X = x)

If highly successful

220000

0.5

110000

If moderately successful

130000

0.3

39000

If unsuccessful

0

0.2

.0

Expected net profit = r * P(X = 1)

Expected net profit = 149000

Profit probability ** P(X = 1)
220000 0.5 24200000000
130000 0.3 5070000000
0 0.2 0

Variance = ??*P(X = r) - (E(x))

= 29270000000 - 149000

= 7069000000

SD = Var

= 84077.3453

(b)

20% opportunity cost of the investment of $300000 = 20% * 300000

= $ 60000

This is less than the expected net profit in (a) which was $149000.

Therefore he should undertake the project

(c) This can be done similarly as the (a)

I am using absolute values for easy calculations

Return(%) Pobability x* P(X=x)
10 0.3 3
22 0.2 4.4
20 0.5 10

Expected % return = r * P(X = 1)

Expected % return = 17.4%

Variance = ??*P(X = r) - (E(x))   

= 326.8 - 17.4^{2}

= 24.04 % %

SD = Var

= 4.903%